# Information Inequalities for Five Random Variables

### v1

@@ 7.1. Case I
For notational convenience $\mathsfit b\mkern 1.0mu(x,y)$, for binomial, denotes the function defined on $\mathbb{N}\times\mathbb{N}$ that satisfies the following recurrent definition for positive integers $x$ and $y$:
$\mathsfit b\mkern 1.0mu(0,0)$ $=\mathsfit b\mkern 1.0mu(x,0)=\mathsfit b\mkern 1.0mu(0,y)=1,\mbox{ and}$
$\mathsfit b\mkern 1.0mu(x,y)$ $=\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{-}\mkern 1.5mu1,y)+\mathsfit b\mkern 1.0mu(x,y\mkern 1.5mu{-}\mkern 1.5mu1).$
Clearly, $\mathsfit b\mkern 1.0mu(x,y)=\mathsfit b\mkern 1.0mu(y,x)={x+y\choose x}$. The following summation formulas will be used later.
@@
$\sum_{i\leqslant x,\;j\leqslant y}\mathsfit b\mkern 1.0mu(i,j)$ $=\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)-1,$
$\sum_{i\leqslant x,\;j\leqslant y}i\mkern 1.5mu\mathsfit b\mkern 1.0mu(i,j)$ $=x\mkern 1.5mu\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)-\mathsfit b\mkern 1.0mu(x,y\mkern 1.5mu{+}\mkern 1.5mu1)+1.$
@@
Induction on $x$ shows that $\sum_{i\leqslant x}\mathsfit b\mkern 1.0mu(i,y)=\mathsfit b\mkern 1.0mu(x,y\mkern 1.5mu{+}\mkern 1.5mu1)$, and also that
${\textstyle\sum}_{i\leqslant x}i\mkern 1.5mu\mathsfit b\mkern 1.0mu(i,y)=x\mkern 1.5mu\mathsfit b\mkern 1.0mu(x,y\mkern 1.5mu{+}\mkern 1.5mu1)-\mathsfit b\mkern 1.0mu(x\mkern 1.5mu{-}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu2).$
@@
For a finite, downward closed set $S\subset\mathbb{N}\times\mathbb{N}$ define the $3$-dimensional vector $\mathbf{v}_{S}$ as
$\mathbf{v}_{S}=\langle\alpha_{S},\beta_{S},\gamma_{S}\rangle=\sum_{(i,j)\in S}\mathsfit b\mkern 1.0mu(i,j)\,\langle 1,i,j\rangle,$
@@
The diagonal $D_{1}$ has a single point, the origin, and the corresponding vector is $\mathbf{v}_{D_{1}}=\langle 1,0,0\rangle$. In general, the diagonal $D_{n}$ has $n(n+1)/2$ points, and the vector associated with $D_{n}$ is
@@
The following theorem provides a family of non-Shannon inequalities that covers all inequalities that were found experimentally in Section 6, Case I.
@@
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha_{S}\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta_{S}\mkern 1.5mu\mathcal{C}+\gamma_{S}\mathcal{D}\geqslant 0.$ (10)
@@
By (i) it suffices to show that $f^{*}$ satisfies all inequalities in (10). By (ii), permutationally equivalent subsets of $N^{*}$ have the same $f^{*}$-value. Below, $c^{k}$ etc. stands for $k$ elements chosen from $c_{1},\ldots,c_{n}$. Occasionally, $c_{1}$, $d_{1}$ will also be denoted by $c$ and $d$, and $c^{k+1}$ will be written as $cc^{k}$, letting $c=c_{1}$ be one of the chosen elements.
@@
${-(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}d^{\ell}z)}+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k+1}d^{\ell}z)+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}d^{\ell+1}z).$
@@
$\sum_{(k,\ell)\in S}\mathsfit b\mkern 1.0mu(k,\ell)\,\mathcal{I}(k,\ell).$
On the left hand side we have $\alpha_{S}$ many copies of $abcd\mkern 1.5mu]$ and $\mathcal{Z}$, $\beta_{S}$ many copies of $\mathcal{C}$, and $\gamma_{S}$ copies of $\mathcal{D}$. On the right hand side the only remaining negative term is $(a,b\mkern 1.0mu{|}\mkern 1.0muz)$, all others cancel out as $\mathsfit b\mkern 1.0mu(k\mkern 1.5mu{-}\mkern 1.5mu1,\ell)+\mathsfit b\mkern 1.0mu(k,\ell\mkern 1.5mu{-}\mkern 1.5mu1)=\mathsfit b\mkern 1.0mu(k,\ell)$. Consequently, inequality (10) holds in $f^{*}$, as claimed. ∎
@@
${-(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}d^{\ell}z)}+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k+1}d^{\ell}z)+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}d^{\ell+1}z).$
Write the right hand side as
$\big(c^{k}d^{\ell}z-c^{k+1}d^{\ell}z-c^{k}d^{\ell+1}z\big)+{}$ (11)
$\big(abc^{k}d^{\ell}z-abc^{k+1}d^{\ell}z-abc^{k}d^{\ell+1}z\big)+{}$ (12)
$\big(ac^{k+1}d^{\ell}z-ac^{k}d^{\ell}z+ac^{k}d^{\ell+1}z\big)+{}$ (13)
$\big(bc^{k+1}d^{\ell}z-bc^{k}d^{\ell}z+bc^{k}d^{\ell+1}z\big).$ (14)
@@
[…]
The sum of the right hand sides in the estimates (15), (16), (17) and (18) is
@@ 7.2. Case II
Let $f$ be a polymatroid on abcdzabcdz that has an $n$-copy for the partition $\{cd\}\{ab\}\{z\}$. Then $f$ satisfies the following inequality for every $k\leqslant n$:
$(a,b\mkern 1.0mu{|}\mkern 1.0muz)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{Z}+\frac{k(k-1)}{2}\mathcal{C}\geqslant 0.$
@@
$acbd\mkern 1.5mu]+\mathcal{Z}+k\mkern 1.5muC\geqslant-(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}z)+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k+1}z).$ (19)
@@
$\mathcal{C}\geqslant(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k}z)-2\mkern 1.5mu(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k+1}z)+(a,b\mkern 1.0mu{|}\mkern 1.0muc^{k+2}z),$
required by the induction does not hold in general. Instead, we give a more involved reasoning, resembling the proof of Lemma 15. In (19) write $c^{k+1}$ as $cc^{k}$, and let $d$ be the element that forms a pair with this $c$. Adding $(b,d\mkern 1.0mu{|}\mkern 1.0muc^{k}z)+(a,c\mkern 1.0mu{|}\mkern 1.0muc^{k}dz)$ to the right hand side of (19) and rearranging, we obtain the upper bound
@@
(20) and (21) follow from submodularity. (22) expresses that $c$ and $c^{k}z$ are independent over $ab$. (23) and (24) are in Lemma 12, while (25) and (26) are from Lemma 14. The sum of the right hand sides of (20)–(26) is
$acbd\mkern 1.5mu]+\mathcal{Z}+k\mkern 1.5mu\mathcal{C}-(c,z\mkern 1.0mu{|}\mkern 1.0mub)-(d,z\mkern 1.0mu{|}\mkern 1.0mua),$
@@
To describe the technique used in the second proof of Theorem 16, let $\mathcal{E}_{n}$ denote the collection of all linear $5$-variable inequalities that are valid in every polymatroid on abcdzabcdz that has an $n$-copy. Let $f^{*}$ be such an $n$-copy. Having $n$ instances of $cd$, one of the $c_{i}d_{i}$ pairs can be singled out, and one of its elements can be renamed $z^{\prime}$. Restricting $f^{*}$ to these elements is an $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy of $abcdz^{\prime}$ since $z^{\prime}$ and the remaining n1n-1 pairs are independent over $ab$. Therefore, $abcdz^{\prime}$ satisfies the inequalities in $\mathcal{E}_{n-1}$. Let $E(a,b,c,d,z)\in\mathcal{E}_{n-1}$ be such an inequality, marking the base elements explicitly. Then we have $E(a,b,c,d,c)\in\mathcal{E}_{n}$, and also $E(a,b,c,d,d)\in\mathcal{E}_{n}$. This fact has been observed and used in Section 6 to explain the coefficients in the obtained inequalities that do not contain the variable $z$.
@@
$\mathcal{J}(\alpha,\beta)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,b\mkern 1.0mu{|}\mkern 1.0muz)+\alpha\mkern 1.5mu[acbd\mkern 1.5mu]+\alpha\mathcal{Z}+\beta\mkern 1.5mu\mathcal{C}\geqslant 0.$
@@
$(a,b|z)+[acbd\mkern 1.5mu]+\mathcal{Z}$ $\geqslant(az,bz\mkern 1.0mu{|}\mkern 1.0mucz)-3(cd,z\mkern 1.0mu{|}\mkern 1.0muab),$
$[acbd\mkern 1.5mu]+\mathcal{Z}$ $\geqslant[az,cz,bz,dz]-3(cd,z\mkern 1.0mu{|}\mkern 1.0muab),$
$(a,c\mkern 1.0mu{|}\mkern 1.0mub)$ $\geqslant(az,cz\mkern 1.0mu{|}\mkern 1.0mubz)-(c,z\mkern 1.0mu{|}\mkern 1.0muab),$
$(b,c\mkern 1.0mu{|}\mkern 1.0mua)$ $\geqslant(bz,cz\mkern 1.0mu{|}\mkern 1.0muaz)-(c,z\mkern 1.0mu{|}\mkern 1.0muab).$
Since $cd$ and $z$ are conditionally independent over $ab$ in $f^{*}$, the last terms are zero. Taking the first inequality once, the second one $\alpha$ times, and the last two $(\alpha+\beta)$ times, the sum of the left-hand sides is $\mathcal{J}(\alpha\mkern 1.5mu{+}\mkern 1.5mu1,\alpha\mkern 1.5mu{+}\mkern 1.5mu\beta)$, while the right hand side is just $\mathcal{J}(\alpha,\beta)$ for the $(az,bz,cz,dz,cz)$ base. Since this inequality is in $\mathcal{E}_{n-1}$ by assumption, we have $\mathcal{J}(\alpha\mkern 1.5mu{+}\mkern 1.5mu1,\alpha\mkern 1.5mu{+}\mkern 1.5mu\beta)\in\mathcal{E}_{n}$, as claimed. ∎
Proof II of Theorem 16.
@@
Lemma 17 remains valid if in the definition of $\mathcal{J}(\alpha,\beta)$ the Ingleton expression $acbd\mkern 1.5mu]$ is replaced by $abcd\mkern 1.5mu]$. Consequently, some, but not all, of the inequalities covered in Theorem 11 can be obtained by similar inductive reasoning.

### v2

@@ 7.1. Case I
For notational convenience $\mathbfit b(x,y)$, for binomial, denotes the function defined on $\mathbb{N}\times\mathbb{N}$ that satisfies the following recurrent definition for positive integers $x$ and $y$:
$\mathbfit b(0,0)$ $=\mathbfit b(x,0)=\mathbfit b(0,y)=1,\mbox{ and}$
$\mathbfit b(x,y)$ $=\mathbfit b(x\mkern 1.5mu{-}\mkern 1.5mu1,y)+\mathbfit b(x,y\mkern 1.5mu{-}\mkern 1.5mu1).$
Clearly, $\mathbfit b(x,y)=\mathbfit b(y,x)={x+y\choose x}$. Table 4 illustrates the initial values of $\mathbfit b(x,y)$.
$\scriptstyle 0$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 0$$\scriptstyle 1$$\scriptstyle 2$$\scriptstyle 3$$\scriptstyle 4$$\scriptstyle 5$$\scriptstyle 6$$\scriptstyle 7$1144101020203535565684841201201133661010151521212828363611223344556677881111111111111111yyxx
Table 4. The function $\mathbfit b(x,y)$
The following summation formulas will be used later.
@@
$\sum_{i\leqslant x,\;j\leqslant y}\mathbfit b(i,j)$ $=\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)-1,$
$\sum_{i\leqslant x,\;j\leqslant y}i\mkern 1.5mu\mathbfit b(i,j)$ $=x\mkern 1.5mu\mathbfit b(x\mkern 1.5mu{+}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu1)-\mathbfit b(x,y\mkern 1.5mu{+}\mkern 1.5mu2)+1.$
@@
Induction on $x$ shows that $\sum_{i\leqslant x}\mathbfit b(i,y)=\mathbfit b(x,y\mkern 1.5mu{+}\mkern 1.5mu1)$, and also that
${\textstyle\sum}_{i\leqslant x}i\mkern 1.5mu\mathbfit b(i,y)=x\mkern 1.5mu\mathbfit b(x,y\mkern 1.5mu{+}\mkern 1.5mu1)-\mathbfit b(x\mkern 1.5mu{-}\mkern 1.5mu1,y\mkern 1.5mu{+}\mkern 1.5mu2).$
@@
For a finite, downward closed set $s\subset\mathbb{N}\times\mathbb{N}$, define the three-dimensional vector $\mathbf{v}_{s}$ as
$\mathbf{v}_{s}=\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle=\sum_{(i,j)\in s}\mathbfit b(i,j)\,\langle 1,i,j\rangle,$
@@
For illustration, the $13$ downward-closed subsets of the diagonal $D_{3}$, as well as the corresponding $\langle\alpha_{s},\beta_{s},\gamma_{s}\rangle$ triplets are presented in Table 5.
$\scriptstyle\#1$$\scriptstyle\#2$$\scriptstyle\#3$$\scriptstyle\#4$$\scriptstyle\#5$$\scriptstyle\#6$$\scriptstyle\#7$$\scriptstyle\#8$$\scriptstyle\#9$$\scriptstyle\#10$$\scriptstyle\#11$$\scriptstyle\#12$$\scriptstyle\#13$
#1\#1 $\langle 1,0,0\rangle$
#2\#2 $\langle 2,1,0\rangle$
#3\#3 $\langle 3,3,0\rangle$
#4\#4 $\langle 2,0,1\rangle$
#5\#5 $\langle 3,1,1\rangle$
#6\#6 $\langle 4,3,1\rangle$
#7\#7 $\langle 5,3,3\rangle$
#8\#8 $\langle 6,5,3\rangle$
#9\#9 $\langle 3,0,3\rangle$
#10\#10 $\langle 4,1,3\rangle$
#11\#11 $\langle 5,3,3\rangle$
#12\#12 $\langle 6,3,5\rangle$
#13\#13 $\langle 7,5,5\rangle$
Table 5. Downward closed subsets of $D_{3}$ and the corresponding triplets
For example, the set #6\#6 has elements $(0,0)$, $(1,0)$, $(2,0)$, and $(0,1)$; the associated triplet is
$\mathbfit b(0,0)\langle 1,0,0\rangle+\mathbfit b(1,0)\langle 1,1,0\rangle+\mathbfit b(2,0)\langle 1,2,0\rangle+\mathbfit b(0,1)\langle 1,0,1\rangle=\langle 4,3,1\rangle.$
The diagonal $D_{1}$, which is #1\#1 in the list, has a single point, the origin, and the corresponding vector is $\mathbf{v}_{D_{1}}=\langle 1,0,0\rangle$. In general, the diagonal $D_{n}$ has $n(n+1)/2$ points, and the vector associated with $D_{n}$ is
@@
The following theorem provides a family of non-Shannon inequalities that covers all inequalities that were found experimentally in Section 6.1.
@@
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+\alpha_{s}\big([abcd\mkern 1.5mu]+\mathcal{Z}\big)+\beta_{s}\mkern 1.5mu\mathcal{C}+\gamma_{s}\mathcal{D}\geqslant 0.$ (11)
@@
By (i) it suffices to show that $f^{*}$ satisfies all inequalities in (11). By (ii), permutationally equivalent subsets of $N^{*}$ have the same $f^{*}$-value. Below, $c^{k}$, etc., stands for $k$ elements chosen from $c_{1},\ldots,c_{n}$. Occasionally, $c_{1}$, $d_{1}$ will also be denoted by $c$ and $d$, and $c^{k+1}$ will be written as $cc^{k}$, letting $c=c_{1}$ be one of the chosen elements.
@@
${-(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}d^{\ell}z)}+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k+1}d^{\ell}z)+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}d^{\ell+1}z).$
@@
$\sum_{(k,\ell)\in s}\mathbfit b(k,\ell)\,\mathcal{I}(k,\ell).$
On the left hand side we have $\alpha_{s}$ many copies of $abcd\mkern 1.5mu]$ and $\mathcal{Z}$, $\beta_{s}$ many copies of $\mathcal{C}$, and $\gamma_{s}$ copies of $\mathcal{D}$. On the right hand side the only remaining negative term is $(a,b\mkern 1.5mu{|}\mkern 1.5muz)$, all others cancel out as $\mathbfit b(k\mkern 1.5mu{-}\mkern 1.5mu1,\ell)+\mathbfit b(k,\ell\mkern 1.5mu{-}\mkern 1.5mu1)=\mathbfit b(k,\ell)$. Consequently, inequality (11) holds in $f^{*}$, as claimed. ∎
@@
${-(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}d^{\ell}z)}+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k+1}d^{\ell}z)+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}d^{\ell+1}z).$
Write the right hand side as the sum $T_{1}+T_{2}+T_{3}+T_{4}$ where the four terms are
$T_{1}$ $=c^{k}d^{\ell}z-c^{k+1}d^{\ell}z-c^{k}d^{\ell+1}z,$ (12)
$T_{2}$ $=abc^{k}d^{\ell}z-abc^{k+1}d^{\ell}z-abc^{k}d^{\ell+1}z,$ (13)
$T_{3}$ $=ac^{k+1}d^{\ell}z-ac^{k}d^{\ell}z+ac^{k}d^{\ell+1}z,$ (14)
$T_{4}$ $=bc^{k+1}d^{\ell}z-bc^{k}d^{\ell}z+bc^{k}d^{\ell+1}z.$ (15)
@@
$T_{4}\leqslant bcz+k(bc-b)+(\ell+1)(bd-b).$ (19)
The sum of the right hand sides in the estimates (16)–(19) is
@@ 7.2. Case II
Let $f$ be a polymatroid on abcdzabcdz that has an $n$-copy for the partition $\{cd\}\{ab\}\{z\}$. Then $f$ satisfies the following inequality for every $1\leqslant k\leqslant n$:
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+k\mkern 1.5mu[acbd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{Z}+\frac{k(k-1)}{2}\mathcal{C}\geqslant 0.$
@@
$acbd\mkern 1.5mu]+\mathcal{Z}+k\mkern 1.5mu\mathcal{C}\geqslant-(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}z)+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k+1}z).$ (20)
@@
$\mathcal{C}\geqslant(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k}z)-2\mkern 1.5mu(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k+1}z)+(a,b\mkern 1.5mu{|}\mkern 1.5muc^{k+2}z),$
required by the induction does not hold in general. Instead, we give a more involved reasoning, resembling the proof of Lemma 15. In (20) write $c^{k+1}$ as $cc^{k}$, and let $d$ be the element that forms a pair with this $c$. Adding $(b,d\mkern 1.5mu{|}\mkern 1.5muc^{k}z)+(a,c\mkern 1.5mu{|}\mkern 1.5muc^{k}dz)$ to the right-hand side of (20) and rearranging, we obtain the upper bound
@@
Equations (21) and (22) follow from submodularity. Equation (23) expresses that $c$ and $c^{k}z$ are independent over $ab$. Equations (24) and (25) are in Lemma 12, while (26) and (27) are from Lemma 14. The sum of the right-hand sides of (21)–(27) is
$acbd\mkern 1.5mu]+\mathcal{Z}+k\mkern 1.5mu\mathcal{C}-(c,z\mkern 1.5mu{|}\mkern 1.5mub)-(d,z\mkern 1.5mu{|}\mkern 1.5mua),$
@@
To describe the technique used in the second proof of Theorem 16, let $\mathcal{E}_{n}$ denote the collection of all linear five-variable inequalities that are valid in every polymatroid on abcdzabcdz that has an $n$-copy. Let $f^{*}$ be such an $n$-copy. Having $n$ instances of $cd$, one of the $c_{i}d_{i}$ pairs can be singled out, and one of its elements can be renamed $z^{\prime}$. Restricting $f^{*}$ to these elements is an $n\mkern 1.5mu{-}\mkern 1.5mu1$-copy of $abcdz^{\prime}$ since $z^{\prime}$ and the remaining n1n-1 pairs are independent over $ab$. Therefore, $abcdz^{\prime}$ satisfies the inequalities in $\mathcal{E}_{n-1}$. Let $E(a,b,c,d,z)\in\mathcal{E}_{n-1}$ be such an inequality, marking the base elements explicitly. Then we have $E(a,b,c,d,c)\in\mathcal{E}_{n}$, and also $E(a,b,c,d,d)\in\mathcal{E}_{n}$. This fact has been observed and used in Section 6 to explain the coefficients in the obtained inequalities that do not contain the variable $z$.
@@
$\mathcal{J}(\alpha,\beta)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(a,b\mkern 1.5mu{|}\mkern 1.5muz)+\alpha\mkern 1.5mu[acbd\mkern 1.5mu]+\alpha\mathcal{Z}+\beta\mkern 1.5mu\mathcal{C}\geqslant 0.$
@@
$(a,b|z)+[acbd\mkern 1.5mu]+\mathcal{Z}$ $\geqslant(az,bz\mkern 1.5mu{|}\mkern 1.5mucz)-3(cd,z\mkern 1.5mu{|}\mkern 1.5muab),$
$[acbd\mkern 1.5mu]+\mathcal{Z}$ $\geqslant[az,cz,bz,dz]-3(cd,z\mkern 1.5mu{|}\mkern 1.5muab),$
$(a,c\mkern 1.5mu{|}\mkern 1.5mub)$ $\geqslant(az,cz\mkern 1.5mu{|}\mkern 1.5mubz)-(c,z\mkern 1.5mu{|}\mkern 1.5muab),$
$(b,c\mkern 1.5mu{|}\mkern 1.5mua)$ $\geqslant(bz,cz\mkern 1.5mu{|}\mkern 1.5muaz)-(c,z\mkern 1.5mu{|}\mkern 1.5muab).$
Since $cd$ and $z$ are conditionally independent over $ab$ in $f^{*}$, the last terms are zero. Taking the first inequality once, the second one $\alpha$ times, and the last two $(\alpha+\beta)$ times, the sum of the left-hand sides is $\mathcal{J}(\alpha\mkern 1.5mu{+}\mkern 1.5mu1,\alpha\mkern 1.5mu{+}\mkern 1.5mu\beta)$, while the right-hand side is just $\mathcal{J}(\alpha,\beta)$ for the $(az,bz,cz,dz,cz)$ base. Since this inequality is in $\mathcal{E}_{n-1}$ by assumption, we have $\mathcal{J}(\alpha\mkern 1.5mu{+}\mkern 1.5mu1,\alpha\mkern 1.5mu{+}\mkern 1.5mu\beta)\in\mathcal{E}_{n}$, as claimed. ∎
Proof 2 of Theorem 16.
@@
Lemma 17 remains valid if, in the definition of $\mathcal{J}(\alpha,\beta)$, the Ingleton expression $acbd\mkern 1.5mu]$ is replaced by $abcd\mkern 1.5mu]$. Consequently, the inequality
$(a,b\mkern 1.5mu{|}\mkern 1.5muz)+k\mkern 1.5mu[abcd\mkern 1.5mu]+k\mkern 1.5mu\mathcal{Z}+\frac{k(k-1)}{2}\mathcal{C}\geqslant 0$
is also in $\mathcal{E}_{n}$ for all $k\leqslant n$. This inequality is covered by Theorem 11 when the downward closed set $s$ consists of the lattice points $(0,i)$ for $0\leqslant i\leqslant k$.